OpenStudy (anonymous):

A team has won 60% of the games it has played so far. If the team wins the 15 remaining games in the season, it will have won 80% of the season's games. How many games will the team play over the entire season?

OpenStudy (anonymous):

Let t be the total of games played and w be the number of games won at the point in time where the team has 15 games to play in the season. Solve the following simutaneous equations for t and w.\[\frac{w}{t}=\frac{60}{100},\frac{w+15}{t+15}= \frac{80}{100}\]\[\{w\to 9,t\to 15\} \]The sum of 15 + 15 = 30 is the total numbers of games over the entire season.

OpenStudy (anonymous):

Sorry. "numbers of games" should have been typed "number of games".